# X-GAN: A Generative AI-Powered Unsupervised Model for High-Precision Segmentation of Retinal Main Vessels toward Early Detection of Glaucoma

### v1

@@ Abstract
Structural changes in main retinal blood vessels serve as critical biomarkers for the onset and progression of glaucoma. Identifying these vessels is vital for vascular modeling yet highly challenging. This paper proposes X-GAN, a generative AI-powered unsupervised segmentation model designed for extracting main blood vessels from Optical Coherence Tomography Angiography (OCTA) images. The process begins with the Space Colonization Algorithm (SCA) to rapidly generate a skeleton of vessels, featuring their radii. By synergistically integrating generative adversarial networks (GANs) with biostatistical modeling of vessel radii, X-GAN enables a fast reconstruction of both 2D and 3D representations of the vessels. Based on this reconstruction, X-GAN achieves nearly 100% segmentation accuracy without relying on labeled data or high-performance computing resources. Also, to address the Issue, data scarity, we introduce GSS-RetVein, a high-definition mixed 2D and 3D glaucoma retinal dataset. GSS-RetVein provides a rigorous benchmark due to its exceptionally clear capillary structures, introducing controlled noise for testing model robustness. Its 2D images feature sharp capillary boundaries, while its 3D component enhances vascular reconstruction and blood flow prediction, supporting glaucoma progression simulations. Experimental results confirm GSS-RetVein’s superiority in evaluating main vessel segmentation compared to existing datasets. Code and dataset are here: https://github.com/VikiXie/SatMar8.
@@ 1 Introduction
Glaucoma is one of the primary causes of irreversible blindness and vision impairment [35, 37, 58]. Early detection is crucial, as the disease often progresses silently, with noticeable symptoms appearing only after significant damage has already occurred [21, 22]. Recent studies have identified changes in choroidal blood vessels, especially microvasculature dropout, as promising new biomarkers for monitoring glaucoma progression [41, 54]. These changes can be assessed using the advanced imaging technology of Optical Coherence Tomography Angiography (OCTA) techniques [30, 23]. However, accurately evaluating choroidal vessel density using OCTA is a complex and nontrivial task.
Existing medical imaging analytics in glaucoma research, such as extraction and segmentation of major retinal blood vessels, predominantly depend on computer vision techniques [29, 27, 31, 28, 23, 37, 58, 57, 38, 36, 35, 19]. In recent years, supervised learning segmentation models (SLSMs) [49, 19, 44, 42, 31, 27, 28, 16, 39, 29, 62, 6, 64, 9] have become particularly prominent in this field. However, SLSMs may not be well-suited for detecting choroidal vessels in OCTA images. Supervised learning approaches rely heavily on labeled data, but manually annotating choroidal blood vessels in OCTA images is extremely challenging with current devices [30, 40]. OCTA images offer high-resolution, depth-resolved visualization of retinal and choroidal microvasculature without the need of contrast dye, making them more effective than traditional fundus imaging for detecting vascular abnormalities [30, 40]. However, segmenting choroidal vessels is difficult due to irregular patterns, intersecting pathways, dense capillary networks [26, 21], and complex capillary structures. Moreover, the presence of major blood vessels, which must be excluded, further complicates the labeling process [21, 22].
To fill the gap, this research introduces a novel unsupervised approach to detecting blood vessels in OCTA images.
Glaucoma researchers have revealed that retinal blood vessels follow biostatistical relationships, meaning that vessel radius variations along branching structures [7, 23]. raising the question of whether segmentation can bypass pixel-based mapping and leverage vessel radius as a key marker. To address this, we propose X-GAN, an unsupervised model that integrates GANs [13] with retinal vessel biostatistics for precise main vessel segmentation from OCTA images, eliminating the need for labeled data or extensive training. To tackle data scarcity, we introduce GSS-RetVein, a benchmark dataset with clear capillary distribution in 2D and 3D, enhancing segmentation, 3D reconstruction, and blood flow prediction.
@@
  • We introduce X-GAN, a generative AI-powered unsupervised segmentation model tailored for extracting main blood vessels from OCTA images.
  • We propose the GSS-RetVein, a unique and pioneering dataset focused exclusively on retinal vascularity. GSS-RetVein consists of 2D and 3D imaging data, and also preserves intricate capillary details without compromising resolution, ensuring sharp blood vessel boundaries and minimal pixel-induced blurring.
  • Experimental results of X-GAN on GSS-RetVein, OCTA-500 [30] and ROSE [40] for retinal main vessel segmentation have demonstrate that X-GAN outperformed state-of-the-art (SOTA) cureent models, achieving nearly 100% segmentation accuracy. Also, experimental results of different models on GSS-RetVein have shown the superiority of GSS-RetVein.
@@ 2 Related Work
2.1 Datasets for Glaucoma
Many datasets of glaucoma focus on computer vision tasks [37, 58, 57, 38, 36, 35, 28, 40, 30, 63, 52, 53]. For inner retinal blood vessels, datasets like LAG [28], OCTA-500 [30] and ROSE [40], provide various-sizeimages (e.g., 3mm, 6mm). However, these images represent only a small portion of the total dataset. Some datasets [63, 52, 53, 37, 58, 57, 38, 36, 35, 28] offer fundus images, and the capillaries are nearly indistinguishable to the naked eye. There is a shortage of high-quality, intuitive data, such as detailed images of inner 3mm×$\times$3mm retinal vessels, and existing datasets often lack sufficient clarity. Consequently, few studies focus on structural changes in major retinal blood vessels in glaucoma.
Figure 1: The architecture of X-GAN. Red indicates that X-GAN does not utilize these things in Segmentor $S$, whereas GAN+SLSMs do.
2.2 Computer Vision for Glaucoma
Even when data is accessible, challenges in labeling and the subjective variability among doctors may introduce biases or hinder optimal model fitting [37, 58, 57, 38, 36, 35, 19]. Generative AI can compensate for data deficiencies and labeling issues, though it often requires human verification [60] and generated data also face challenges in labeling. However, GAN [13, 66, 20, 65, 24, 61] is an unsupervised model which can solve both issues: data scarity and labeling. GAN-based segmentation methods are generally designed in conjunction with a dedicated segmentation network. For the segmentor part $S$, many pre-trained models like CNNs [27, 31, 28], U-Net [49, 64, 19, 44], ViT [10], Mask R-CNN [16] and MedSAM [39] can be used for final segmentation. Most recent studies from CVPR [50, 20, 2, 8], IEEE Transactions on Medical Imaging [23, 61, 11, 24] and other top conferences or journals [59, 3, 65] also follow this way.
@@ 2.3 Motivation
For GAN+SLSM [23, 61, 2, 11, 24, 20], or SLSMs [49, 19, 44, 42, 31, 27, 28, 16, 39, 29, 62, 6, 64, 9], pixel-level representations can be subjective, as different doctors may annotate the details of the same object differently [35, 30, 29]. In OCTA imaging, retinal blood vessels exhibit a uniform color change and have a shape closely resembling a path. They can be represented using coordinates along with road width. Space Colonization Algorithm (SCA) [46] just meets this requirement. Rather than generating pixel-wise vessel segmentations, we formulates the vascular network as a structured graph representation, ensuring topological consistency and robust connectivity. The graph consists of the following components, shown in Eq. 1:
@@
We introduce the Depth-First Search (DFS, a nineteenth century algorithm)-based segmentation approach [55], applying DFS directly to structured (coordinates, radius) data before image rendering. Unlike GAN+SLSM [23, 61, 2, 11, 24, 20], which segments generated images mixed with real data, our method intercepts the GAN generation process, ensuring segmentation aligns with biostatistical vessel radii. The DFS threshold, $R_{min}$, is defined as the minimum radius ratio of main vessels, following established biostatistical principles [22, 12, 5, 43, 7].
Also, we introduce GSS-RetVein, a dataset capturing rich details of both main vessels and capillaries. As shown in Fig. 4, its images (2D, Fig. 4(b)) retain clear capillary edges and tips, unlike ROSE [40] (Fig. 4(d)) and OCTA-500 [30] (Fig. 4(c)). The dense capillaries introduce noise, effectively testing model robustness for SLSMs. This detail aids vascular reconstruction, blood flow prediction, and glaucoma progression simulation by enhancing edge and endpoint learning [30, 29, 62]. Inspired by the Harvard glaucoma dataset series [37, 58, 57, 38, 36, 35], we integrate fairness measures to minimize bias and ensure equitable model performance.
Note: given the model’s simplicity (adjusted GAN + DFS) yet crazy performance (nearly 100% segmentation accuracy), we name it ”X” to highlight its exceptional segmentation ability, similar to how Elon Musk rebranded Twitter as X. Thus, we refer to our model as X-GAN.
3 Methodology: X-GAN
3.1 Vessel Structure Refinement Module
Given a vessel structure $X_{s}$ via SCA, our generator $G$ learns to transform it into a realistic OCTA-like representation $X_{r}^{\prime}$, expressed as $G(X_{s})$, while $X_{r}^{\prime}$ undergoes style adaptation to align with the contrast and noise characteristics of real OCTA images $X_{r}$. To enhance structural consistency, we modify the original CycleGAN architecture by removing the inverse generator and incorporating a segmentation consistency loss via the segmentor $S$. This modification prevents CycleGAN from altering the vessel topology, ensuring that vessel segmentation remains biologically meaningful, as illustrated in Eq. 2:
@@ 3.1 Vessel Structure Refinement Module
where Eq. 3 is the Wasserstein Adversarial Loss, Eq. 4 is the Gradient Penalty for Stability ($\hat{X}$ is the reference vessel representation) and Eq. 5 is the Structural Consistency via Segmentation Loss. Upon training convergence, the generator produces a vessel structure that not only exhibits realistic contrast properties but also preserves anatomical coherence, ensuring fidelity to real OCTA characteristics.
@@ 3.2 Segmentor
First, we construct a graph $G=(V,E)$, where nodes $V$ represent vessel centerline points and edges $E$ connect adjacent vessel points based on local vessel connectivity. A vessel segment $e_{ij}$ between nodes $v_{i}$ and $v_{j}$ (the direction is from $i$ to $j$) is assigned a weight, as shown in Eq. 6:
@@
where larger vessels are prioritized in the traversal process. To extract primary vessels, we apply radius thresholding and DFS traversal. For radius filtering, we define a minimum main vessel radius $R_{min}$ and retain only Eq. 7:
@@
For DFS traversal, it consists of three steps: (1) select the optic disc region as the root node; (2) recursively traverse the largest connected component using DFS; (3) stop when no further vessel segments satisfy the radius constraint. When all three steps finished, construct final extracted vessel structure $G_{main}$. This approach effectively retains only the primary vascular structure, removing small capillaries and noise. This method can be summarized as the following Algorithm 1.
Algorithm 1 : Primary Vessel Extraction via DFS
1:Vessel structure $V=\{(x_{i},y_{i},z_{i}|r_{i})\}_{i=1}^{N}$
2:Extracted primary vessel graph $G_{\text{main}}$
3:Step 1: Vessel Graph Construction
4:for all vessel point $(x_{i},y_{i},z_{i}|r_{i})\in V$ do
5:     Identify neighboring vessel points to form edges $E$
6:     Assign edge weights: $w(e_{ij})=r_{i}$
7:end for
8:Step 2: Vessel Filtering
9:Apply radius threshold $R_{\text{min}}$:
10:$V_{\text{main}}=\{(x_{i},y_{i},z_{i}|r_{i})\mid r_{i}\geq(R_{\text{min}}*r_{max})\}$
11:Remove weakly connected components
12:Step 3: DFS for Vessel Traversal
13:Initialize DFS from optic disc region
14:for all vessel point $(x_{i},y_{i},z_{i}|r_{i})\in V_{\text{main}}$ do
15:     Recursively explore vessel network using DFS
16:     Expand along larger vessels
17:end for
18:Step 4: Output Primary Vessel Graph return $G_{\text{main}}$
The key aspect of Algorithm 1 is the threshold $R_{min}$. Instead of the final image, we analyze intermediate outputs $S(X_{r})$, CSV data (Fig. 2(e)), and GAN-derived radius statistics (Fig. 2(d)). For a sample image (Fig. 2(a)), we computed a Mean (85.31) and Standard Deviation (88.67), determining $R_{min}$ as 0.1968 $R_{max}$. Across datasets, values were 0.2011 (GSS-RetVein), 0.2102 (OCTA-500), and 0.1982 (ROSE), standardized to 0.2 for consistency. Ophthalmology studies [22, 12, 5, 43, 7] confirm 18-26% capillary coverage in OCTA images, validating this threshold.
(a) GSS-RetVein
(b) 3D Visualization
(c) Statistical Distribution
(d) Radius Statistics
(e) Vascular Data Format (Part)
Figure 2: 3D visualization of the pixel distribution of one sample, and its Gaussian statistical distribution.
[…]
4 Dataset and Implement
@@ 4.1 Dataset
4.1.1 GSS-RetVein
GSS-RetVein is a unique and pioneering dataset focused exclusively on retinal vascularity. The dataset features an exceptionally clear distribution of capillaries. It comprises 550 images, both 2D and 3D, from 250 subjects, ensuring fair representation across gender and race. We collected retinal vascular data for both eyes (OS and OD) of each subject, with a gender split of 48.8% male and 51.2% female. Racial distribution reflects local glaucoma clinic demographics: 80.4% white, 14.6% Black, with Asians and other minority groups comprising the remainder. Additional data includes age (47.1 $\pm$ 24.8 years) and examination dates from 2023 to 2024.
@@ 4.1.1 GSS-RetVein
GSS-RetVein’s original images come from the latest Intalight device [^1] https://intalight.com/, offering superior detail compared to OPTOVUE [^2] https://www.visionix.com/ (Fig. 3). As shown in Fig. 3(a) and Fig. 3(c), Intalight images feature clearer vessel boundaries and capillary distributions, enhancing segmentation accuracy. In supervised learning, their minimal pixel variation introduces valuable noise for testing model robustness [62]. For 3D reconstruction and evolution prediction, these high-resolution images aid in vascular modeling and blood flow analysis [29], making Intalight data ideal for both tasks.
(a) Intalight
(b) OPTOVUE
(c) Intalight
(d) OPTOVUE
Figure 3: Comparison of data collected by different devices for the same patient. Among them, Fig. 3(a) and Fig. 3(b) are images of deep angio, and Fig. 3(c) and Fig. 3(d) are images of superficial angio. They all are images of inner retina vessel.
For GSS-RetVein, we eliminate red and green baselines from Intalight device images (Fig. 4(a)) through a multi-step process. First, we apply a dark channel algorithm [15] to remove light artifacts, followed by unsharp masking [51] to enhance high-frequency details. The final images (Fig. 4(b)) undergo manual verification for quality assurance. As shown in Fig. 4, while main vessels remain visible across datasets, magnification reveals that OCTA-500 [30] and ROSE [40] blur capillaries and vessel ends. Even the unprocessed baseline image (Fig. 4(a)) offers superior vessel clarity compared to other datasets (Fig. 4(d), Fig. 4(c)).
(a) Baseline
(b) GSS-RetVein
(c) OCTA-500
(d) ROSE
Figure 4: Comparison of GSS-RetVein images with other datasets.
For the representation of 3D imaging of GSS-RetVein, as shown in Fig. 5, capillaries stand out from the thicker main blood vessels by attaching to them and narrowing at their ends (Fig. 5(c)). The elaborate architecture of blood vessels and the intertwining capillaries contribute to an extremely intricate vascular layout. 3D imaging part of GSS-RetVein is formatted as a JSON file, which includes 3D X, Y, and Z coordinates and the corresponding radius pixel distribution values. For annotation (2D, image format: PNG), our team includes two glaucoma specialists who handle labeling, verification, and manual evaluation, utilizing the Labelme [^3] https://github.com/wkentaro/labelme.
(a) Top View
(b) Front View
(c) Detail
Figure 5: 3D imaging of retinal vessel shown from various angles.
In general, GSS-RetVein is a dataset that combines 2D and 3D images with extremely high image resolution, and is particularly dedicated to retinal vascular segmentation.

### v3

# X-GAN: A Generative AI-Powered Unsupervised Model for Main Vessel Segmentation of Glaucoma Screening

@@ Abstract
Structural changes in main retinal blood vessels serve as critical biomarkers for the onset and progression of glaucoma. Identifying these vessels is vital for vascular modeling yet highly challenging. This paper proposes X-GAN, a generative AI-powered unsupervised segmentation model designed for extracting main blood vessels from Optical Coherence Tomography Angiography (OCTA) images. The process begins with the Space Colonization Algorithm (SCA) to rapidly generate a skeleton of vessels, featuring their radii. By synergistically integrating the generative adversarial network (GAN) with biostatistical modeling of vessel radii, X-GAN enables a fast reconstruction of both 2D and 3D representations of the vessels. Based on this reconstruction, X-GAN achieves nearly 100% segmentation accuracy without relying on labeled data or high-performance computing resources. Experimental results confirm X-GAN’s superiority in evaluating main vessel segmentation compared to existing deep learning models. Code is here: https://github.com/VikiXie/SatMar8.
Keywords:
Medical Imaging Unsupervised Learning Glaucoma Pathological Segmentation Generative AI
@@ 1 Introduction
Glaucoma is a leading cause of irreversible blindness, often progressing silently until significant vision loss occurs [40, 37, 38, 42, 30]. Recent research highlights choroidal microvasculature dropout as a potential biomarker for disease progression [42]. Optical Coherence Tomography Angiography (OCTA) enables visualization of these vascular changes, but accurately assessing choroidal vessel density remains a challenging task [1, 17].
Existing medical imaging analytics in glaucoma research predominantly depend on computer vision techniques [2, 45, 44, 43, 17, 37, 38, 36, 39, 41, 40, 46]. In recent years, supervised learning segmentation models (SLSMs) [10, 46, 12, 13, 44, 45, 43, 9, 8, 2, 3, 14, 7, 15] have become particularly prominent in this field. However, SLSMs may not be well-suited for detecting choroidal vessels in OCTA images. They rely heavily on labeled data, but manually annotating choroidal blood vessels in OCTA images is extremely challenging with current devices [1, 4]. OCTA images offer high-resolution, depth-resolved visualization of retinal and choroidal microvasculature without the need of contrast dye, making them more effective than traditional fundus imaging for detecting vascular abnormalities [1, 4]. However, segmenting choroidal vessels is difficult due to irregular patterns, intersecting pathways, dense capillary networks [42], and complex capillary structures. Moreover, the presence of major blood vessels, which must be excluded, further complicates the labeling process [42, 30].
Glaucoma researchers have revealed that retinal blood vessels follow biostatistical relationships, meaning that vessel radius variations along branching structures [32, 17]. raising the question of whether segmentation can bypass pixel-based mapping and leverage vessel radius as a key marker. To address this, we propose X-GAN, an unsupervised model that integrates GANs [34] with retinal vessel biostatistics for precise main vessel segmentation from OCTA images, eliminating the need for labeled data or extensive training.
@@
  • We introduce X-GAN, a generative AI-powered unsupervised segmentation model tailored for extracting main blood vessels from OCTA images without annotations.
  • We eliminate the reliance on synthetic label supervision mechanisms (SLSMs) by introducing a novel segmentation strategy based on vessel radius thresholds and depth-first search (DFS), enabling accurate main vessel extraction without any pixel-level annotations.
  • Experimental results of X-GAN for retinal main vessel segmentation have demonstrate that X-GAN outperformed state-of-the-art (SOTA) current models, achieving nearly 100% segmentation accuracy.
@@ 2 Related Work
2.1 Medical Imaging in Glaucoma
Many datasets of glaucoma focus on computer vision tasks [37, 38, 36, 39, 41, 40, 43, 4, 1, 53, 54, 52], especially in medical image processing. For inner retinal blood vessels, datasets like LAG [43], OCTA-500 [1] and ROSE [4], provide various-sizeimages (e.g., 3mm, 6mm). However, these images represent only a small portion of the total dataset. Some datasets [53, 54, 52, 37, 38, 36, 39, 41, 40, 43] offer fundus images, and the capillaries are nearly indistinguishable to the naked eye. There is a shortage of high-quality, intuitive data, such as detailed images of inner 3mm×$\times$3mm retinal vessels, and existing datasets often lack sufficient clarity. Consequently, few studies focus on structural changes in major retinal blood vessels in glaucoma.
2.2 AI for Glaucoma
Even with accessible data, labeling remains challenging due to annotation costs and inter-observer variability, introducing bias and hindering model performance [37, 38, 36, 39, 41, 40, 46]. Generative AI can alleviate data scarcity and labeling constraints, though human oversight is often required [50]. GANs [34, 35, 18, 24, 27, 25], as unsupervised models, address both issues by generating realistic samples without manual labels. In this framework, segmentation is typically achieved by coupling the GAN with a dedicated segmentor $S$, which can be implemented using pre-trained CNNs [45, 44, 43], U-Net variants [10, 7, 46, 12], ViT [22], Mask R-CNN [9], or MedSAM [8], enabling end-to-end vessel synthesis and segmentation.
@@ 2.3 Motivation
In traditional models such as GAN+SLSM [17, 25, 20, 26, 27, 18] and standalone SLSMs [10, 46, 12, 13, 44, 45, 43, 9, 8, 2, 3, 14, 7, 15], pixel-level segmentations are prone to inter-observer variability [40, 1, 2]. However, in OCTA, retinal vessels have consistent visual traits, uniform grayscale gradients and path-like shapes, making them suitable for coordinate-based modeling with associated widths.
We adopt the Space Colonization Algorithm (SCA) [49] to represent vessels as structured graphs, enhancing topological coherence and connectivity. To bridge the domain gap between initial vessel maps and real OCTA data, we apply a CycleGAN-based refinement module [35], aligning generated structures with realistic image characteristics while preserving topology. Unlike prior methods that segment post-rendered images, we introduce a DFS-based segmentation approach [51] applied directly to graph-structured (coordinates, radius) data. This intercepts the generation process to ensure anatomical fidelity, governed by a radius threshold $R_{min}$ derived from biostatistical vessel metrics [30, 31, 33, 32].
3 Methodology
Figure 1: The architecture of X-GAN where red indicates that X-GAN does not utilize these things in Segmentor $S$, whereas GAN+SLSMs do.
3.1 Vessel Structure Refinement Module
Rather than generating pixel-wise vessel segmentations, we formulates the vascular network as a structured graph representation, ensuring topological consistency and robust connectivity. The graph consists of the following components, shown in Equation 1:
@@ 3.1 Vessel Structure Refinement Module
Given a vessel structure $X_{s}$ via SCA, our generator $G$ learns to transform it into a realistic OCTA-like representation $X_{r}^{\prime}$, expressed as $G(X_{s})$, while $X_{r}^{\prime}$ undergoes style adaptation to align with the contrast and noise characteristics of real OCTA images $X_{r}$. To enhance structural consistency, we modify the original CycleGAN architecture by removing the inverse generator and incorporating a segmentation consistency loss via the segmentor $S$. This modification prevents CycleGAN from altering the vessel topology, ensuring that vessel segmentation remains biologically meaningful, as illustrated in Equation 2:
@@
where Equation 3 is the Wasserstein Adversarial Loss, Equation 4 is the Gradient Penalty for Stability ($\hat{X}$ is the reference vessel representation) and Equation 5 is the Structural Consistency via Segmentation Loss. Upon training convergence, the generator produces a vessel structure that not only exhibits realistic contrast properties but also preserves anatomical coherence, ensuring fidelity to real OCTA characteristics.
@@ 3.2 Segmentor
First, we construct a graph $G=(V,E)$, where nodes $V$ represent vessel centerline points and edges $E$ connect adjacent vessel points based on local vessel connectivity. A vessel segment $e_{ij}$ between nodes $v_{i}$ and $v_{j}$ (the direction is from $i$ to $j$) is assigned a weight, as shown in Equation 6:
@@
where larger vessels are prioritized in the traversal process. To extract primary vessels, we apply radius thresholding and DFS traversal. For radius filtering, we define the ratio of the smallest main vessel radius to the largest main vessel radius $R_{min}$ and retain only Equation 7:
@@
For DFS traversal, it consists of three steps: (1) select the optic disc region as the root node; (2) recursively traverse the largest connected component using DFS; (3) stop when no further vessel segments satisfy the radius constraint. When all three steps finished, construct final extracted vessel structure $G_{main}$. This approach effectively retains only the primary vascular structure, removing small capillaries and noise.
4 Dataset and Implementation
@@ 4.1 Dataset
4.1.1 OCTA-500
4.1.2 ROSE
(a) UTSW Data
(b) OCTA-500
(c) ROSE
Figure 2: Comparison of UTSW with Other Datasets
4.1.3 UTSW
our partner medical institutions’s institutional review board (IRB) approved this study, which followed the principles of the Declaration of Helsinki. Since the study was retrospective, the IRB waived the requirement for informed consent from patients.
@@ 4.1.3 UTSW
As show in Fig. 2, original images come from the latest Intalight device [^1] https://intalight.com/, which is currently the most advanced equipment. As shown in Fig. 2(a), Intalight images feature clearer vessel boundaries and capillary distributions, enhancing segmentation accuracy. UTSW comprises 550 images, both 2D and 3D, from 250 subjects, ensuring fair representation across gender and race. We collected retinal vascular data for both eyes (OS and OD) of each subject, with a gender split of 48.8% male and 51.2% female. Racial distribution reflects local glaucoma clinic demographics: 80.4% white, 14.6% Black, with Asians and other minority groups comprising the remainder. Additional data includes age (47.1 $\pm$ 24.8 years) and examination dates from 2023 to 2024. In supervised learning, their minimal pixel variation introduces valuable noise for testing model robustness [3]. For 3D reconstruction and evolution prediction, these high-resolution images aid in vascular modeling and blood flow analysis [2], making Intalight data ideal for both tasks.
Our team includes two attending glaucoma physicians who are responsible for the detailed labeling of the main retinal blood vessels using the tool Labelme [^2] https://github.com/wkentaro/labelme.
4.2 Implementation
4.2.1 Hyperparameter
For X-GAN, its generator adopts a ResNet 9-block architecture, while the discriminator utilizes PatchGAN (70$\times$70). The optimizer is Adam with a learning rate of $2\times 10^{-4}$ and momentum parameters $\beta_{1}=0.5$, $\beta_{2}=0.999$. X-GAN is trained for 50 epochs, with a linear learning rate decay to 0 after 25 epochs, with evaluation performed using 10-fold cross-validation. Data augmentation includes: random rotations ($k\times 90^{\circ}\pm 10^{\circ}$) and flipping.
4.2.2 Evaluation Matrix
We use Intersection over Union (IoU), also known as the Jaccard Index, to measure similarity between the predicted segmentation and the ground truth. This metric ranges from 0 to 1, with 1 indicating perfect segmentation. Additionally, we use the Dice Coefficient, another metric for segmentation accuracy that calculates overlap between the prediction and ground truth. All in all, IoU is used to evaluate segmentation accuracy of the main vessels, while the Dice Coefficient assesses accuracy at finer vessel ends relative to the main vessels. For vascular structural enhancement, we use Structural Similarity Index (SSIM) and Mean Squared Error (MSE). SSIM (range: -1 to 1) assesses similarity based on luminance, contrast, and structure, while MSE quantifies pixel-wise differences, with lower values indicating higher similarity. While SSIM aligns with human perception, MSE focus on absolute pixel differences.
4.2.3 Experimental Setup
We selected the following deep learning models as reference benchmarks: U-Net [10], U-Net++ [7], Attention U-Net [12], Mask R-CNN [9], YOLOv11-x [11], MedSAM [8], CauSSL [13], UniverSeg [14], S2VNet [15] and Tyche [16]. These models are leaders in the fields of image segmentation and medical image processing. For their parameter adjustment, we use Optuna [^3] https://optuna.org/ to optimize their learning green, decay and other hyperparameters. For the training and testing of the model, we use ten cross-validation methods. The evaluation indicators are finally averaged.
4.2.4 Hardware
The hardware for training and testing include 2 Tesla V100 GPUs (2 $\times$ 32GB), 64GB of RAM, 8 CPU cores per node, and a total of 6 nodes.
5 Experimental Result
5.1 Comparative Experiment
Table 1: Comparison X-GAN with Other Models on Different Datasets ($\times$100%)